### v3

# ResBit: Residual Bit Vectors for Categorical Values

Abstract
One-hot vectors, a method for representing discrete/categorical data, are commonly used in machine learning due to their simplicity and intuitiveness. However, the one-hot vectors suffer from a linear increase in dimensionality, posing computational and memory challenges, especially when dealing with datasets containing numerous categories. To address this issue, we propose Residual Bit Vectors (ResBit), a technique for densely representing categorical data. While Analog Bits (Chen et al., 2023) presents a similar approach, it faces challenges in categorical data generation tasks. ResBit overcomes these limitations, offering a more versatile solution. In our experiments, we focus on tabular data generation, examining the performance across scenarios with varying amounts of categorical data. We verify the acceleration and ensure the maintenance or improvement of performance.
Machine Learning, ICML, tabular data generation, curse of dimensionality
1 Introduction
Figure 1: Overview of the proposed ResBit (left) and application example to TabDDPM (right).
The application of machine learning in real-world scenarios is increasing, and one notable case is the use of tabular data. Tabular data consists of complex information, encompassing numerical data representing continuous values and categorical data representing discrete quantities. Consequently, decision tree-based methods without using deep learning still dominate tabular data analysis due to its intricate nature (Grinsztajn et al., 2022). Handling categorical data in machine learning involves converting it into numerical form, and deep learning commonly employs one-hot encoding for this purpose. Examples of this encoding technique extend beyond tabular data analysis and are observed in other domains, such as label or class assignment in image classification and class-conditional image generation.
One-hot vectors are widely utilized due to their simplicity and ease of implementation. However, they come with drawbacks such as high memory consumption due to sparsity and an increase in computational complexity as dimensions grow. When considering the representation of categorical data using one-hot vectors, the challenges posed by the “curse of dimensionality” make it difficult to handle in machine learning. The issue arises from the increased dimensionality leading to increased memory usage and computational demands, making the practical implementation in machine learning challenging.
In this paper, we focus on tabular data generation and consider the solution to the “curse of dimensionality” associated with one-hot vectors. The research on tabular data generation evaluates its performance using various datasets. Each paper provides information about the datasets used, including not only the data size but also details such as the type of task and the number of columns. However, in many cases, despite the conversion of categorical data to one-hot vectors, there has been a lack of examination regarding the increase in dimensionality. Figure 2 presents our investigation into the cardinality of the categorical features for datasets used in recent tabular data generation research. This figure shows the total cardinalities for each data. The per-column cardinality of the each data is shown in Table 10.
Figure 2: Plot of the total cardinalities of each data.
From this figure, it is evident that the cardinality is low for most datasets. When considering the application to real-world scenarios, we take the example of Credit Card Transaction Data. In such data, information about transactions includes details like ‘what was purchased?’ and ‘where it was purchased?’. In typical scenarios with such datasets, the cardinality of categorical data is often extremely high  [^1] For example, the zip code in the U.S. has 41,690 unique values.. Considering the application of machine learning to real-world scenarios, verification in this aspect becomes essential.
To address the dimensionality increase of one-hot vectors accompanying the growth in cardinality, we propose Residual Bit Vectors (ResBit) by integrating the concepts of Analog Bits (Chen et al., 2023) and Residual Vector Quantization (RVQ) (Juang & Gray, 1982) (Section 3.4). ResBit represents categorical data by hierarchically acquiring bit representations. Our research is motivated by the phenomenon of generation collapse observed in TabDDPM (Kotelnikov et al., 2023), one of the tabular data generative models. This phenomenon, akin to mode collapse, involves the generation of only specific values. This problem has also been encountered in other studies using TabDDPM as a baseline, as noted in the work of Zhang et al. (2023). While maintaining the same hyperparameter settings during the training of TabDDPM, we confirmed that the phenomenon of generation collapse is dependent on the cardinality of categorical data. This dependence is observed in conjunction with the dimensionality of the input to the multinomial diffusion models constituting TabDDPM, where the input is provided in the form of one-hot vectors. Based on this observation, we concluded that the dimensionality of one-hot vectors is a contributing factor to generation collapse. While conventional bit representations suffice for representing categorical data, they introduce specific problem in generation tasks (Section 3.3). ResBit addresses this problem while mitigating the increase in dimensionality even with growing cardinality (Section 3.5).
To confirm the effectiveness of ResBit, we conduct experiments by applying it to TabDDPM (see Figure 1), which was the motivation for our research (Section 4). Following the concept of analog bits, treating ResBit as numerical data allowed us to maintain performance in low cardinality datasets while generating diverse categorical values in high cardinality datasets. Additionally, we revealed advantages and disadvantages of existing tabular data generative models that have not been previously discussed. Furthermore, our experimental results contribute to providing answers to the question of how to handle categorical data in tabular data generation.
2 Backgrounds
2.1 Tabular Data Generation with Deep Learning
Generating tabular data using deep learning is highly meaningful. Unlike vision and NLP tasks, tabular data has limited data size and more directly includes personal information compared to text or images. Therefore, using synthetic data rather than raw data is an example of privacy consideration. TVAE and CTGAN (Xu et al., 2019) are models that apply VAE (Kingma & Welling, 2013) and GAN (Goodfellow et al., 2014) to tabular data generation, respectively, and are still used as baselines in the field. In addition to those, there is a distinct approach utilizing diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Research has been conducted on methods such as TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023), STaSy (Kim et al., 2023), and TabSyn (Zhang et al., 2023). These methods have demonstrated high performance compared to GAN-based approaches.
2.2 Cardinality of the Categorical Features
Figure 2 shows the total cardinality of the each dataset used in recent tabular data generation researches. For more details, please refer to the Appendix B. Many datasets shown in Figure 2 have small cardinality for categorical features, making it challenging to assess the suitability of each method for real-world applications. For example, in e-commerce sites, an item ID may have high cardinality, and it is not demonstrated whether generation models can effectively handle such cases. IDs like these often represent entirely different entities with slight variations, making it difficult to treat them simply as continuous values. In this study, experiments are conducted on high cardinality datasets to assess the performance of existing methods.
2.3 Discrete/Categorical Data in Diffusion Models
The issue mentioned in Section 2.2 is similar to the treatment of tokens in text generation. Tokens, unlike semantic vectors, may represent entirely different meanings with slight variations in value. Consequently, research has been conducted on employing methods that handle discrete values, such as categorical data or text, with diffusion models originally designed for continuous values. D3PM (Austin et al., 2021) and Multinomial Diffusion (Hoogeboom et al., 2021) are approaches that convert the noise distribution, initially assumed to be a normal distribution for continuous values, into a categorical distribution. Although the input is provided as a one-hot vector, these methods can learn imbalanced distributions as they predict categorical distributions. These techniques require designing models to adapt to categorical distributions. Bit Diffusion (Chen et al., 2023) represents discrete data in binary and models analog bits, treated as continuous values, using Gaussian diffusion. It is characterized by the high expressiveness of diffusion models, eliminating the need for model designs tailored to discrete data. Diffusion-LM (Li et al., 2022) is a method that learns diffusion models for word vectors, and similar ideas are applied to tabular data. TabSyn (Zhang et al., 2023) projects each column into a latent space and handles discrete quantities by processing them in continuous space. The treatment of categorical data involves two main strategies: treating it as a categorical distribution or converting it into continuous values. It is still unclear which approach is optimal. Our method treats categorical data as continuous values, representing a modification in the way these data are expressed.
3 Proposed Method
In this section, we first discuss the preliminaries and observation of the phenomenon of generation collapse in TabDDPM. Subsequently, we introduce the challenges of applying analog bits to tabular data generation tasks (Section 3.3) and its solution, ResBit (Section 3.4).
3.1 Preliminaries
TabDDPM
(Kotelnikov et al., 2023) is the first study in the world to introduce diffusion models to tabular data generation. The tabular data consists of two types of data: numerical and categorical. In TabDDPM, numerical data is modeled using Gaussian diffusion models, while categorical data is modeled using multinomial diffusion (Hoogeboom et al., 2021). The model is represented by MLP, and the coherence of the two types of data is maintained by integrating the losses. For numerical data, a simplified loss following DDPM (Ho et al., 2020) is used as follows.
$L_{t}^{\mathrm{simple}}=\mathbb{E}_{t,\boldsymbol{x}_{0},\varepsilon~{}\sim\mathcal{N}(0,\boldsymbol{I})}[\|\varepsilon-\varepsilon_{\theta}(\boldsymbol{x}_{t},t)\|_{2}^{2}]$
Here, $\boldsymbol{x}_{0}$ denotes the real data and $\varepsilon_{\theta}$ denotes the neural network that predicts the noise $\varepsilon$ added at time step $t$. The categorical data also adopts the loss used in multinomial diffusion models. The predefined distributions for categorical data are represented as $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})=\mathrm{Cat}(\boldsymbol{x}_{t};(1-\beta_{t})\boldsymbol{x}_{t-1}+\beta_{t}/K)$ when the number of classes is $K$. The loss functions for both types are combined to train the loss function of TabDDPM.
$L_{t}^{\mathrm{TabDDPM}}=L_{t}^{\mathrm{simple}}+\dfrac{\sum_{i\leq C}L_{t}^{i}}{C}$
$L_{t}^{i}$ represents two KL divergences between $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})$ and $p_{\theta}(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t})$ for each column.
Analog Bits
is a preprocessing method proposed for handling discrete data in Bit Diffusion (Chen et al., 2023). Normally, discrete data is represented as a one-hot vector, but analog bits represents it as bit strings. Furthermore, by treating bit strings as numerical data, continuous state diffusion models can be applied to discrete data without devising a specific architecture for discrete data.
3.2 Generation Collapse in TabDDPM
Table 1 shows whether a mode collapse-like phenomenon occurred when we randomly sampled 55,000 cases from IBM’s synthetic credit card data [^2] https://ibm.ent.box.com/v/tabformer-data and changed the number of unique values for the data of each category. The percentage here is the percentage of the total training data, and data with frequencies lower than this percentage are masked [^3] This is represented as cat_min_frequency in the official TabDDPM implementation.. Although the dataset used here is only a partial selection, it contains a very large number of categorical values. In this experiment, the hyperparameters except the value indicating the masking ratio are fixed, so changes in this ratio are linked to changes in the dimensionality of the one-hot vector given as input to the multinomial diffusion models. On the basis of this experiment, we hypothesize that the increase in the dimensionality of the one-hot vector affects the success or failure of training.
Table 1: Wether collapse does occur or not when the percentage of masking was changed in the experiment
Threshold (Rate) collapsed?
0.00038 No
0.00028 Yes
0.00019 Yes
0.00010 Yes
For a theoretical explanation of this phenomenon, please refer to Appendix C.
3.3 Out of Index
In binary representation, there is a possibility of generating values that do not exist in the generation task. As an example of this problem, let’s consider representing the states in the United States. There are 50 states in the United States, and if we represent them using one-hot vectors, it would be like $\boldsymbol{x}\in\{0,1\}^{50}$. Specifically, for $j$-th state, $x_{j}=1$ and $x_{i}=0$ for $i\neq j$. Representing this in binary would be possible with a 6-dimensional representation. The model generates each element as either 0 or 1, resulting in $642^{6}=$ 64 possible combinations. However, out of these combinations, 14 would represent state numbers that do not correspond to any existing states. For example, if the model predicts $(1,1,0,1,0,1)$, it corresponds to the number 53, but this value cannot be mapped to any existing state in the context of representing the 50 states of the United States. We conveniently refer to this problem as out of index, and an overview is provided in Figure 3. Increasing the number of trainings could potentially address this problem. However, this approach is effective when the training data is extensive, and for tabular data generation, which involves training on individual datasets, there is a simple risk of overfitting.
Figure 3: Overview of the “out of index” problem in the states of the U.S. example.
3.4 Residual Bit Vectors
A straightforward solution to address the “out of index” problem discussed in Section 3.3 is to limit categorical data to one special string and $12^{n}-$ 1 different categories. However, this method imposes limitations on the inherent diversity of the training data. Other approaches such as clipping or normalizing out-of-range values are also conceivable. However, the clipping method introduces issues where the distribution of generated data may change. In the case of normalization, as mentioned in Section 2.3, there is a risk that slight value differences might refer to entirely different entities, making this solution highly inappropriate. In essence, the maximum representable number should be the number of categorical data. In the example of U.S. states mentioned earlier, as there are 50 states, a representation method that indicates integers from 0 to 49 inclusive would be necessary.
We propose Residual Bit Vectors (ResBit) inspired by RVQ to fulfill the aforementioned requirement. RVQ recursively applies vector quantization to the quantization error obtained when quantizing the original vector. ResBit accomplishes this using a binary representation. Figure 4 provides an overview of ResBit. When using ResBit, like one-hot vectors, prior knowledge of the number of classes is required. Let $M$ represent the number of classes. There exists a non-negative integer $b_{1}$ such that inequality (1) is satisfied. Representing $12^{b_{1}}-$ 1 in binary results in a bit sequence where all elements are 1s, and its length is $b_{1}$.
$12^{b_{1}}-1\leq M-1<2^{b_{1}+1}-$ 1 (1)
In this state, it can represent values from 0 to $12^{b_{1}}-$ 1, but cannot represent values outside this range. Therefore, the same process is applied to $M-(2^{b_{1}}-1)$. This repetition is how the representation of ResBit is obtained. Formally, the required ResBit for representing $M$ classes, indexed from 0, can be written as follows:
$M-1=\sum_{k=0}^{\infty}a_{k}(2^{k}-1)\quad\mathrm{s.t.}\min\sum_{k=0}^{\infty}a_{k}k$ (2)
The $k$-th term represents integers in the range $[0,a_{k}(2^{k}-1)]$. An integer $N\in\mathbb{Z}$, where $0\leq N<M$, can be represented as follows:
$N=\sum_{k=1}^{\infty}c_{k}(2^{k}-1)\quad(c_{k}\in[0,a_{k}],c_{k}\in\mathbb{Z})$ (3)
Finally, we present the example of U.S. states mentioned earlier. Since $M=50$, we have $350-1=49=1\cdot(2^{5}-1)+1\cdot(2^{4}-1)+1\cdot(2^{2}-1)=31+15+$ 3. In other words, U.S. states can be represented in $11(=5+4+2)$ dimensions. Using this approach, we can represent all integers from 0 to 49 inclusive. For example, for the case of 39, which is $39=31+8$, the ResBit representation is $39=((1,1,1,1,1),(1,0,0,0),(0,0))=(1,1,1,1,1,1,0,0,0,0,0)$. This example is illustrated in Figure 4. Pseudo-code for ResBit is provided in Appendix A.7.
Figure 4: ResBit in the states of the U.S. example.
3.5 Decreasing the Dimensionality
Due to its properties, ResBit can reduce the increase in dimensionality compared to one-hot vectors. Figure 5 illustrates the comparison of the required dimensionality for representing integers less than $M$ when the number of classes is $M$. For example, when $M=1,000$, the required dimensionality is 1,000 for one-hot vectors, while ResBit only requires 42 dimensions, resulting in a reduction of 95.8%. This allows us to alleviate the “curse of dimensionality” associated with using one-hot vectors.
Figure 5: Comparison of the dimensionality between one-hot and ResBit when the number of classes are increased to $10^{6}$.
4 Experiments
We consider Table Residual Bit Diffusion (TRBD), which is incorporated ResBit into TabDDPM (see Figure 1). In this section, we demonstrate the effectiveness of integrating ResBit into TabDDPM through experiments. When introducing ResBit into TabDDPM, it is challenging to directly apply ResBit as multinomial diffusion assumes one-hot vectors as inputs. Therefore, following the concept of analog bits, ResBit is treated as numerical data. ResBit is combined with numerical data, and preprocessing is performed using QuantileTransformer. Specifically, One-Hot Encoder is replaced by Residual Bit Encoder, and its outputs are treated as numerical data following the idea of analog bits. After converting each categorical value to its ResBit, the process described in Equation 4 is performed. This is the same process as that of TabDDPM.
$x_{cat_{i}}^{resbit}=\log(\max(\texttt{ResBit\_Encoder}(x_{cat_{i}}),10^{-30}))$ (4)
Here, ResBit_Encoder converts the category value to ResBit, and $x_{cat_{i}}$ is the $i$-th categorical column. After converting categorical values to ResBit, we combine them with numerical data and further transform them using Quantile Transformer.
$x_{in}=\mathrm{QuantileTransformer}(\mathrm{concat}(x_{num},x_{cat_{1}}^{resbit},\ldots))$
4.1 Experimental Setups
We briefly describe the baselines, datasets, and evaluation methods. For software information and further details, please refer to Appendices A.1A.2, and A.4.
Baselines.
We use six existing methods as baselines. The first two are CTGAN (Xu et al., 2019) and TVAE (Xu et al., 2019) as the de fact standards. Additionally, we include TabDDPM (Kotelnikov et al., 2023), CoDi  (Lee et al., 2023), StaSy (Kim et al., 2023), and TabSyn (Zhang et al., 2023) as the diffusion-based methods. These methods were nearly simultaneously introduced, and comprehensive comparisons have not been extensively conducted. Zhang et al. (2023) conducted a comparison, but the evaluation was based on only six datasets, which is limited compared to 11 datasets (Lee et al., 2023) or 15 datasets (Kotelnikov et al., 2023; Kim et al., 2023).
Datasets.
We use 10 datasets to conduct a comprehensive evaluation of the existing methods and verify the effectiveness of our proposed method. The basic overview of the dataset is presented in Table 2. For additional details such as the number of training data, please refer to Table 9 in Appendix A.4. ResBit is a method representing categorical data. Therefore, experiments are conducted on datasets that include both categorical and numerical data. To validate the effectiveness of our proposed method, we prepare datasets with both low and high cardinalities. Low cardinality datasets primarily involve datasets used in TabDDPM (Kotelnikov et al., 2023). There are two main purposes for evaluating on datasets of these two different scales. One purpose is to verify the performance of the proposed method on low cardinality datasets. The other is to assess the effectiveness of the proposed method and existing methods on high cardinality datasets. Most datasets are real-world datasets, except for the CC dataset, which is a synthetic dataset.
Table 2: Brief overview of the dataset used in our experiments.
Dataset Name Abbreviation Cardinality
Credit Card (IBM) CC high
Airlines AR
Insurance IS low
Buddy BD
Adult AD
Churn Modeling CH
Cardio CA
King KI
Abalone AB
Facebook Comm. Vol. FB
Evaluation methods
We evaluate the existing and our methods from three aspects. 1) cardinality of the categorical features: compare the number of generated categories in each categorical column, aiming for a metric that reflects proximity to the training data. 2) TSTR (Train in Synthesis, Test on Real) (Esteban et al., 2017; Yoon et al., 2019): train a classifier on generated data and evaluate its performance on test data to measure sampling quality. We use CatBoost (Prokhorenkova et al., 2018) as the classifier, and conduct five measurements, reporting the average and standard deviation. Experiments are conducted with 10,000 samples for the low cardinality datasets and 3,000,000 samples for the high cardinality datasets. 3) runtime: confirm the impact of applying ResBit on the runtime.
4.2 TSTR Framework
We show the results in Tables 3 and 4. Additional results can be found in Appendix D. Failed indicates that the evaluation using TSTR was not possible due to insufficient diversity in the generated labels for the classification task. In addition, if the training did not complete within 24 hours, the result is marked as N/A, adhering to the criteria outlined in the generative trilemma (Xiao et al., 2022). The generative trilemma outlines speed, quality, and diversity as the three requirements for the sampling stage in generative tasks. However, we argue that these criteria should also be applicable during the training phase. The ideal scenario for a generative task is to achieve diverse and high-quality generation rapidly through short training. Therefore, we believe that investing time in training solely to meet the requirements of the generative trilemma is inappropriate.
Table 3: Results of TSTR framework in classification datasets. Metric is AUROC.
CC AR BD AD CH CA
Real .971$\pm$.009 .727$\pm$.004 .931$\pm$.001 .927$\pm$.002 .855$\pm$.010 .805$\pm$.000
TVAE Failed .640$\pm$.002 .978$\pm$.001 .865$\pm$.004 .763$\pm$.012 .641$\pm$.010
CTGAN Failed .639$\pm$.009 .955$\pm$.004 .874$\pm$.002 .674$\pm$.015 .749$\pm$.007
TabDDPM .376$\pm$.075 .528$\pm$.013 .984$\pm$.001 .909$\pm$.003 .873$\pm$.002 .803$\pm$.001
CoDi N/A N/A .472$\pm$.039 .495$\pm$.062 .584$\pm$.017 .497$\pm$.034
STaSy N/A N/A .978$\pm$.002 .900$\pm$.002 .835$\pm$.004 .791$\pm$.001
TabSyn .930$\pm$.012 .690$\pm$.001 .984$\pm$.001 .899$\pm$.003 .839$\pm$.003 .791$\pm$.001
TRBD (ours) .746$\pm$.051 .638$\pm$.003 .983$\pm$.001 .906$\pm$.000 .870$\pm$.002 .803$\pm$.001
Table 4: Results of TSTR framework in regression datasets. Metric is $R^{2}$ score.
IS KI AB FB
Real .861$\pm$.034 .906$\pm$.002 .557$\pm$.002 .838$\pm$.001
TVAE .634$\pm$.026 .650$\pm$.002 .278$\pm$.006 .539$\pm$.009
CTGAN $-0.706$$\pm$.095 .628$\pm$.012 .291$\pm$.029 .420$\pm$.012
TabDDPM .894$\pm$.003 .864$\pm$.112 .567$\pm$.012 .662$\pm$.008
CoDi $-4.506$$\pm$.616 $-88.074$$\pm$20.908 $-2.518$$\pm$.414 N/A
STaSy .618$\pm$.080 .867$\pm$.010 $-0.613$$\pm$.047 .670$\pm$.011
TabSyn .834$\pm$.004 .867$\pm$.005 .522$\pm$.005 .616$\pm$.009
TRBD (ours) .908$\pm$.003 .892$\pm$.004 .560$\pm$.006 .706$\pm$.005
Comparison of the baselines.
We observe that CoDi requires a substantial amount of time for training when the size of the training data increases. The size of the datasets that Lee et al. (2023) used is up to 36k. However, the size of the FB dataset is 157k. Therefore, it can be inferred that the impact of the dataset size on training time could not be identified. As evident from Table 10, the cardinality of the datasets used by Kim et al. (2023) is low, making it difficult to assess the impact of cardinality on training time. However, the experiments reveal that high cardinality datasets indeed require a significant amount of time for training. Indeed, as demonstrated, existing methods may not perform well on high cardinality datasets, an aspect that has not been widely recognized until now.
Roughly, similar results to Zhang et al. (2023) are obtained, but in our experiments, TabDDPM consistently shows superior performance, suggesting the need for further extensive validation. Additionally, in the large-scale datasets (CC and AR), TabSyn demonstrates superiority. This might be attributed to the Transformer component in TabSyn, which is effective with large scale datasets, possibly related to the requirements of Vision Transformer (Dosovitskiy et al., 2021). Transformers are known to adhere to Scaling Laws (Kaplan et al., 2020), making them more effective with larger datasets. On the other hand, CTGAN outperforms in F1 Score on the AR dataset according to Tables 11, 14, and 18. However, CTGAN fails to generate effectively in the highly imbalanced CC dataset. Thus, it is challenging to conclusively state the superiority of GAN-based methods based on these results.
Evaluation of our proposed method.
From Tables 3 and 4, it can be observed that TRBD maintains performance across most datasets compared to TabDDPM. Moreover, considering the results in Tables 11-20, an overall improvement in performance is evident. In particular, significant performance improvement is observed in the high cardinality datasets (CC and AR). This indicates that multinomial diffusion struggles to handle high cardinality datasets effectively. Furthermore, the reduced standard deviations suggest stable and consistent sampling quality.
Overall discussion.
Taking into account Tables 11-22 in addition to Tables 3 and 4, diffusion models, specifically methods treating categorical data as continuous (TRBD, STaSy, and TabSyn), show a tendency for high sampling quality. In particular, methods that handle categorical and numerical data separately, such as TabDDPM, have not shown competitive results on high cardinality datasets. Additionally, CoDi performs poorly overall. In contrast, STaSy, TabSyn, and TRBD, which process categorical data within the same framework as numerical data, demonstrate competitive results on many datasets. TRBD and TabSyn, which employ dimensionality reduction techniques, also excel on high cardinality datasets, showcasing their effectiveness.
4.3 Cardinality of the Categorical Features
Table 5 shows a comparison of the cardinality of the categorical features. We confirm this by generating an equal number of instances with each generative model. Due to datasets with a high number of columns, it is not possible to show the results for each column. Therefore, we present the ratio to the cardinality of the training data, indicating the maximum ratio from five generations. In low cardinality datasets, the results are similar, whereas in high cardinality datasets, the differences are pronounced. If the models can generate diverse samples, this ratio should approach 1. CoDi seems to produce diverse results, but from Tables 3 and 4, it can be argued that the sampling quality is not good. Both CTGAN and TVAE show a loss of diversity. Neither of them can be considered to achieve the generative trilemma. The lack of diversity is also observed in TabSyn. This suggests the potential loss of fine-grained information due to handling data in the latent space.
Table 5: Comparison of the cardinality of the categorical features in the generated samples.
CC AR IS BD AD CH CA KI AB FB Average
TVAE 0.526 0.917 1.000 0.576 0.618 0.778 0.929 1.000 1.000 0.862 0.821
CTGAN 0.369 0.764 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.899 0.903
TabDDPM 0.487 0.832 1.000 0.870 1.000 1.000 1.000 1.000 1.000 0.973 0.916
CoDi N/A N/A 1.000 1.000 1.000 1.000 1.000 1.000 1.000 N/A 1.000+3N/A
STaSy N/A N/A 1.000 0.924 0.990 1.000 1.000 1.000 1.000 0.963 0.985+2N/A
TabSyn 0.910 1.000 1.000 0.815 0.980 1.000 1.000 1.000 1.000 1.000 0.971
TRBD (ours) 1.000 1.000 1.000 0.957 1.000 1.000 1.000 1.000 1.000 1.000 0.996
4.4 Runtime
Table 6 shows the comparison of the runtime. We use the CH dataset and the CC dataset. The experimental setup has the training batch size of 4,096 and the number of generated samples is 26k for the CH dataset and 100k for the CC dataset. Sampling batch size is equal to the number of generated samples. Other values have not been changed from previous experiments in order to maintain the sampling quality. TabSyn’s training time is the summation of VAE’s and denoising network’s training time. Moreover, note that TabSyn’s denoising network is equipped with early stopping, so training ends early.
Table 6: Comparison of the runtime. OOM denotes out of memory with 24GB GPU.
CH CC
Method Training Sampling Training Sampling
TVAE 33.01s 0.07s 1,973s 33.22s
CTGAN 33.06s 0.22s 36,797s 225s
TabDDPM 567s 22.78s 3,494s OOM
CoDi 6,149s 1.61s N/A N/A
STaSy 4,197s 10.40s N/A N/A
TabSyn 2,325s$+$488s 4.34s 27,377s+2,737s 15.09s
TRBD (ours) 264s 13.57s 183s 4.31s
In the CH dataset, TRBD achieves approximately twice the speedup compared to TabDDPM. Moreover, for the CC dataset, it accomplishes a speedup of almost 20 times during the training phase. TabDDPM and TRBD have the same number of training iterations. Therefore, it is suggested that the computational cost of the multinomial diffusion part in TabDDPM becomes a bottleneck. In other words, considering runtime, TRBD is faster than TabDDPM, capable of generating a wide range of data, and achieving comparable or superior quality. STaSy takes considerably more time, especially for low cardinality datasets, when compared to TabDDPM, TabSyn, and TRBD. This becomes even more pronounced for high cardinality datasets. While the sampling is performed rapidly, satisfying the generative trilemma, the training process is time-consuming. Although the generated quality is competitive, the training time serves as a bottleneck in achieving it. TabDDPM using one-hot vectors demands a significant amount of memory, particularly on the CC dataset. In contrast, TRBD does not exhibit such memory-intensive behavior, resulting in reduced computational costs.
5 Conclusion and Future Work
In this paper, we proposed ResBit as a method to address the “out-of-index” problem in generative tasks. ResBit acquires hierarchical bit representations, overcoming the “curse of dimensionality” faced by one-hot vectors, especially in the case of high cardinality categorical features. In our experiments, when applied to TabDDPM, which motivated our research, ResBit achieved a speedup in training and generation while maintaining accuracy. Additionally, it enabled the generation of diverse categorical values. Through comprehensive experiments, we highlighted shortcomings in existing tabular data generation models. Despite achieving high-speed sampling, training times were notably long, and these models did not demonstrate effective results in configurations not explored in previous researches.
In the future, further validation of ResBit is desired. For instance, although we confirmed its effectiveness in image classification and class-conditional image generation tasks in Appendix E, the experiments were not exhaustive, and there are other potential applications, such as image segmentation or algorithms utilizing one-hot vectors. Furthermore, as existing methods for tabular data generation all convert categorical data into one-hot vectors, constructing models with fewer parameters becomes possible by applying ResBit. While we demonstrated its application to TabDDPM in this paper, its effectiveness is likely to extend to other models as well.
6 Impact Statement
Our research holds the potential for significant dimensionality reduction compared to traditional one-hot vectors by utilizing dense vectors, thereby enabling possibilities such as parameter reduction in deep learning and efficient operations.
Appendix A Details of Experimental Setups
A.1 Experimental Environments
Our software and hardware environments are as follows: Ubuntu 18.04, Python 3.10.8, PyTorch 1.13.1, scikit-learn 1.1.2, and CUDA 10.1, AMD EPYC 7502P 32-Core CPU at 2.5GHz, and one Quadro RTX 6000.
A.2 Baselines
We describe the baseline methods used in our experiments. The baseline methods consist of various types of generative models, including gan-based methods, vae-based methods, and diffusion-based methods.
  • CTGAN and TVAE (Xu et al., 2019) [^4] https://github.com/sdv-dev/CTGAN: The same technique has been applied in various studies to both GANs and VAEs, where it serves as a common baseline in research.
  • TabDDPM (Kotelnikov et al., 2023) [^5] https://github.com/yandex-research/tab-ddpm: This study represents the first introduction of diffusion models for tabular data generation, employing separate models for numerical and categorical data. A shared MLP network is utilized, and learning is achieved by separating the losses. Gaussian diffusion models are employed for numerical data, while multinomial diffusion models are used for categorical data.
  • CoDi (Lee et al., 2023) [^6] https://github.com/ChaejeongLee/CoDi/tree/main: Similar to TabDDPM, numerical data is learned using Gaussian diffusion, and categorical data is learned using multinomial diffusion. However, in this case, two separate networks are employed for learning. The relationships between columns are learned by conditioning the output of each model on the other. Additionally, the use of symmetric learning methods is introduced to strengthen connections.
  • STaSy (Kim et al., 2023) [^7] https://github.com/JayoungKim408/STaSy/tree/main: This diffusion-based generative model employs variance exploding, variance preserving, and sub-variance preserving SDEs (Song et al., 2021) to train the distribution of the tabular data. By incorporating self-placed learning, a curriculum learning technique, and fine-tuning, the method stabilizes the quality of generated samples.
  • TabSyn (Zhang et al., 2023) [^8] https://github.com/amazon-science/TabSyn: This is inspired by Latent Diffusion Models (Rombach et al., 2022). It employs a transformer architecture VAE for projecting into the latent space and an MLP for denoising. The training is designed to embed both numerical and categorical data into the latent space while maintaining their associations, enabling the generation of horizontal connections specific to tabular data. Furthermore, performing denoising in the latent space makes it faster compared to existing diffusion-based methods.
We use official implementations respectively without STaSy. Regarding STaSy, we use the reimplementation [^9] https://github.com/amazon-science/TabSyn/tree/main/baselines/stasy provided by Zhang et al. (2023). If the official implementation has code for searching hyperparameters, it is also executed. We employ Optuna (Akiba et al., 2019) for tuning the hyperparameters.
A.3 Hyperparameter Search Spaces
We present the hyperparameter search spaces in Tables 7 and 8. For evaluating sampling quality, we employ RandomForest (Breiman, 2001) and XGBoost (Chen & Guestrin, 2016) as additional classifiers alongside CatBoost. We utilize Optuna for tuning hyperparameters in CatBoost and grid search to identify the optimal hyperparameters for RandomForest and XGBoost.
Table 7: TRBD hyperparameter spaces from Kotelnikov et al. (2023). Only number of tuning trials is changed to 30.
Parameters Distribution
Learning rate LogUniform[0.00001, 0.003]
Batch size Cat{256, 4096}
Diffusion timesteps Cat{100, 1000}
Training iterations Cat{5000, 10000, 20000}
# MLP layers Int{2, 4, 6, 8}
Width of MLP layers Int{128, 256, 512, 1024}
Proportion of samples Float{0.25, 0.5, 1, 2, 4, 8}
Number of tuning trials 30
Table 8: Hyperparameter space of classifier models from Kim et al. (2023), Lee et al.(2023), and Kotelnikov et al. (2023)
Models Parameters Distribution
CatBoost Learning rate LogUniform[0.001, 1]
Depth UniformInt[3, 10]
L2 leaf reg Uniform[0.1, 10.0]
Bagging temperature Uniform[0, 1]
Leaf estimation iterations UniformInt[1, 10]
XGBoost n_estimators Int{10, 50, 100}
min_child_weight Int{1, 10}
max_depth Int{5, 10}
gamma Float{0.0, 1.0}
nthreads -1
RandomForest max_depth Int{8, 16, Inf}
min_samples_split Int{2, 4}
min_samples_leaf Int{1, 3}
n_jobs -1
A.4 Datasets
Table 9 shows the information of the 10 datasets used in our experiments. This table does not contain the information of label column. All datasets include both numerical and categorical columns. Eight of these datasets are used in the experiment of TabDDPM (Kotelnikov et al., 2023). In addition, we add the two datasets which have high cardinality categorical features.
Table 9: Details of the datasets used in our experiments. C and R denote classification and regression, respectively.
Dataset Name Abbreviation #Num #Cat Task #Train #Test #Val Cardinalities
Credit Card (IBM) CC 5 3 C 384,000 120,000 96,000 (3, 7515, 153)
Airlines AR 3 4 C 345,204 107,877 86,302 (18, 6571, 292, 293)
Insurance IS 3 3 R 856 268 214 (2, 2, 4)
Buddy BD 4 5 C 12,053 3,767 3,014 (3, 56, 19, 10, 4)
Adult AD 6 8 C 26,048 16,281 6,513 (9, 16, 7, 15, 6, 5, 2, 42)
Churn Modeling CH 7 4 R 6,400 2,000 1,600 (3, 2, 2, 2)
Cardio CA 5 6 C 44,800 11,200 14,000 (2, 3, 3, 2, 2, 2)
King KI 17 3 R 13,382 4,323 3,458 (2, 5, 5)
Abalone AB 7 1 R 2,672 836 669 (3)
Facebook Comm. Vol. FB 36 18 R 157,638 19,720 19,722 (81, others are 2)
A.5 Preprocess Datasets
We preprocess the datasets listed in Table 9. The datasets, excluding CC and AR, are obtained from the official TabDDPM implementation [^10] https://www.dropbox.com/s/rpckvcs3vx7j605/data.tar?dl=0. Consequently, we provide details on the preprocessing of the CC and AR datasets.
  • CC: We treat Use Chip, Merchant State, and Merchant City as categorical columns, and missing values as one additional category for them. We delete User, Errors?, and Merchant Name columns. We treat other columns as numerical columns and fill missing values with the column’s average. We convert Year, Month, Day, and Time columns into Timestamp. We sample randomly 600k rows.
  • AR: We treat Airline, Flight, AirportFrom, and AirportTo as categorical columns and other columns as numerical columns. We use all rows.
A.6 Dataset Sources
A.7 Implementation of ResBit
We present the implementation details of ResBit. Algorithm 1 provides pseudocode for computing $a_{k}$ in Equation 2. The variable kind_of_cat represents the number of classes. Algorithm 2 is pseudocode for obtaining ResBit for an integer $N$ in the range $[0,M-1]$, where there are $M$ categories.
Algorithm 1 Get $b_{1},b_{2}\ldots$ Algorithm.
def get_length_ResBit(kind_of_cat: int):
# get length of ResBit for unique category size kind_of_cat
res = []
# 0-index
max_num = kind_of_cat - 1
while True:
bits = bin(max_num)[2:]
# in case of M == 2^m
if "0" not in bits:
res.append(len(bits))
break
else:
s = "1" * (len(bits) - 1)
max_num -= int(s, base=2)
res.append(len(s))
return res
Algorithm 2 Get ResBit Algorithm for $N$.
def int_to_ResBit(N, l):
# l is returns of algorithm 1
res = []
for i in range(len(l)):
if N == 0:
res += [0 for _ in range(l[i])]
continue
bits = "1" * l[i]
X = int(bits, base=2)
if X <= N:
res += [int(z) for z in bits]
N -= X
else:
bits_N = bin(N)[2:]
bits_N = "0" * (l[i] - len(bits_N)) + bits_N
res += [int(z) for z in bits_N]
N = 0
return res
Appendix B Details of the cardinalities survey
Table 10 presents our investigation into the cardinality of the categorical features for datasets used in recent tabular data generation research. This table shows the total cardinalities for each data. Only datasets containing both categorical and numerical columns are shown here. Many studies focus on the number of columns and dataset size, with less emphasis on the cardinality of the categorical features. In Addition, some datasets have unknown cardinality when preprocessing details are not provided in official implementations, and such cases are marked as unknown.
Table 10: Details of the datasets used in recent researches. C and R denote classification and regression, respectively.
Dataset Name #Num #Cat Task Cardinalities References
Adult 6 8 C (9,16,7,15,6,5,2,42) Kotelnikov et al. (2023)
6 9 (9,16,7,15,6,5,2,42,2) Zhang et al. (2023)
Default 14 11 C (81, 2, 7, 56, 11, 11, 11, 11, 10, 10, 2) Zhang et al. (2023)
13 11 Kim et al. (2023)
News 46 2 R (6, 1454) Zhang et al. (2023)
45 14 unknown Kim et al. (2023)
Beijing 7 5 R (5, 12, 31, 24, 4) Zhang et al. (2023)
8 6 unknown Kim et al. (2023)
Insurance 3 3 R (2, 2, 4) Kotelnikov et al. (2023)
3 8 (49, 2, 2, 2, 2, 2, 2, 4) Lee et al. (2023)
Obesity 7 10 C (2, 2, 2, 4, 2, 2, 4, 5, 7, unknown) Kim et al. (2023)
8 9 (2, 2, 2, 4, 2, 2, 4, 5, 7) Lee et al. (2023)
Credit 29 1 C (2) Kim et al. (2023)
HTRU 8 1 C (2)
Spambase 57 1 C (2)
Bean 16 1 C (7)
Crowdsource 28 1 C (6)
Robot 24 1 C (4)
Shuttle 9 1 C (7)
Shoppers 10 8 C (10, 8, 13, 9, 20, 3, 2, 2) Zhang et al. (2023), Kim et al. (2023)
Magic 10 1 C (2)
Buddy 4 5 C (3, 56, 19, 10, 4) Kotelnikov et al. (2023)
Cardio 5 6 C (2, 3, 3, 2, 2, 2)
Churn Modeling 7 4 C (3, 2, 2, 2)
Bank 7 10 C (12, 3, 4, 2, 2, 2, 3, 12, 4, 2) Lee et al. (2023)
Heart 4 10 C (41, 2, 4, 2, 3, 2, 3, 5, 3, 2)
Seismic 11 5 C unknown
Stroke 2 8 C (3, 2, 2, 2, 5, 2, 4, 2)
CMC 2 8 C (4, 4, 2, 2, 4, 4, 2, 3)
Customer 5 7 C (3, 2, 5, 2, 2, 8, 4)
Faults 24 4 C (7, unknown)
Absent 12 9 R (28, 5, 2, 4, 2, 2, others are unknown)
Drug 5 1 R (3)
Abalone 7 1 R (3) Kotelnikov et al. (2023)
Facebook Comm. Vol. 36 18 R (81, others are 2)
King 17 3 R (2, 5, 5)
Appendix C Mutlinomial Diffusion Analysis from the Fomulation
We consider mode collapse from the perspective of the multinomial diffusion equation, the forward process of multinomial diffusion is formulated using a categorical distribution as follows:
$q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})\coloneqq\mathrm{Cat}(\boldsymbol{x}_{t};(1-\beta_{t})\boldsymbol{x}_{t-1}+\beta_{t}/K\cdot\boldsymbol{1})$ (5)
$q(\boldsymbol{x}_{T})\coloneqq\mathrm{Cat}(\boldsymbol{x}_{T};1/K\cdot\boldsymbol{1})$ (6)
$q(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})=\mathrm{Cat}(\boldsymbol{x}_{t}|\overline{\alpha}_{t}\boldsymbol{x}_{0}+(1-\overline{\alpha}_{t})/K\cdot\boldsymbol{1})$ (7)
where $\alpha_{t}\coloneqq 1-\beta_{t},\overline{\alpha}_{t}=\prod_{i\leq t}\alpha_{i}$. $\mathrm{Cat}$ denotes the categorical distribution and $K$ denotes the number of classes. $\beta_{t}$ determines the strength of added noise and satisfies $10<\beta_{1}<\dotsi<\beta_{T}<$ 1. From the equations above, the posterior $q(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})$ can be expressed:
$q(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})=\mathrm{Cat}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{\pi}/\sum_{j=1}^{K}\pi_{j}\right)$ (8)
where $\boldsymbol{\pi}=[\alpha_{t}\boldsymbol{x}_{t}+(1-\alpha_{t})/K\cdot\boldsymbol{1}]\odot[\overline{\alpha}_{t-1}\boldsymbol{x}_{0}+(1-\overline{\alpha}_{t-1})/K\cdot\boldsymbol{1}]$.
We consider a high cardinality scenario, that is, $K\to\infty$. Since $\boldsymbol{x}_{0}$ is represented as a one-hot vector, we have $x_{0,k}=1$ and $x_{0,j}=0$ for $j\neq k$.
$\lim_{K\to\infty}\boldsymbol{\pi}=[\alpha_{t}\boldsymbol{x}_{t}]\odot[\overline{\alpha}_{t-1}\boldsymbol{x}_{0}]=(\alpha_{t}x_{t,i}\cdot\overline{\alpha}_{t-1}x_{0,i})_{0\leq i\leq K}=\alpha_{t}x_{t,k}\cdot\overline{\alpha}_{t-1}$ (9)
Therefore,
$\lim_{K\to\infty}\boldsymbol{\pi}/\sum_{j=1}^{K}\pi_{j}=\boldsymbol{x}_{0}$ (10)
Then,
$q(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{x}_{0})=\mathrm{Cat}(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{0})$ (11)
and we are considering a situation where only a specific class appears.
Appendix D Additional Results of Tabular Data Generation
D.1 TSTR Framework
Tables 11 and 12 present the result of TSTR framework using F1 score and RMSE as metrics. The F1 Score registering as 0 in the CC dataset is likely due to the dataset’s inherent imbalance. While high cardinalities could contribute, it would be expected for the AR dataset to display a similar behavior if that were the sole cause. However, the results in Table 11 suggest otherwise. The AR dataset has a balanced $1:1$ dataset, while the CC dataset is highly imbalanced with a ratio of $0.001:0.999$. These observations lead to the inference that extreme imbalance poses a challenge for generative models, resulting in performance degradation.
Table 11: Additional results of our experiments on the classification datasets. We use CatBoost as the classifier.
CC AR BD AD CH CA
Real .684$\pm$.067 .587$\pm$.004 .987$\pm$.000 .716$\pm$.003 .580$\pm$.033 .738$\pm$.001
TVAE Failed .508$\pm$.002 .846$\pm$.006 .622$\pm$.008 .470$\pm$.020 .510$\pm$.019
CTGAN Failed .584$\pm$.012 .820$\pm$.001 .527$\pm$.007 .249$\pm$.029 .701$\pm$.007
TabDDPM .$0$279$\pm$.134 .907$\pm$.003 .673$\pm$.002 .601$\pm$.007 .737$\pm$.001
CoDi N/A N/A .286$\pm$.030 .357$\pm$.046 .023$\pm$.010 .656$\pm$.008
STaSy N/A N/A .843$\pm$.010 .623$\pm$.008 .574$\pm$.013 .729$\pm$.003
TabSyn .293$\pm$.083 .557$\pm$.004 .905$\pm$.008 .639$\pm$.009 .594$\pm$.010 .721$\pm$.002
TRBD (ours) .$0$482$\pm$.010 .901$\pm$.007 .676$\pm$.004 .612$\pm$.009 .737$\pm$.002
Table 12: Additional results of our experiments on the regression datasets. We use CatBoost as the classifier.
IS KI AB FB
Real 4,475.287$\pm$390.028 107,633.528$\pm$1,022.334 2.009$\pm$.004 5.297$\pm$.019
TVAE 7,769.982$\pm$270.354 230,264.317$\pm$505.661 2.581$\pm$.011 9.115$\pm$.056
CTGAN 16,786.700$\pm$467.101 237,360.868$\pm$3,803.299 2.556$\pm$.052 10.224$\pm$.102
TabDDPM 4,050.754$\pm$47.229 129,491.185$\pm$5,426.840 1.986$\pm$.028 7.648$\pm$.084
CoDi 28,657.977$\pm$1,643.204 3,461,545.525$\pm$412,674.821 6.177$\pm$.373 N/A
STaSy 7,930.060$\pm$857.998 141,939.588$\pm$5,090.508 3.857$\pm$.057 7.710$\pm$.133
TabSyn 5,097.570$\pm$59.577 141,801.203$\pm$2,817.380 2.101$\pm$.012 8.309$\pm$.094
TRBD (ours) 3,785.253$\pm$69.238 115,411.676$\pm$1,972.408 2.002$\pm$.013 7.127$\pm$.057
Tables 13 to 16 present the results using XGBoost as the classifier, while Tables 17 to 20 show the results using RandomForest are shown. Similar trends are observed across other classifiers, as seen in the case of CatBoost.
Table 13: Results of our experiments on the classification datasets. We use XGBoost as the classifier and AUROC as the metric.
CC AR BD AD CH CA
Real .887$\pm$.185 .721$\pm$.001 .987$\pm$.001 .916$\pm$.002 .868$\pm$.007 .799$\pm$.003
TVAE .500$\pm$.000 .638$\pm$.001 .980$\pm$.001 .869$\pm$.002 .760$\pm$.006 .645$\pm$.007
CTGAN .500$\pm$.000 .647$\pm$.003 .962$\pm$.003 .877$\pm$.003 .683$\pm$.009 .758$\pm$.003
TabDDPM .429$\pm$.138 .514$\pm$.007 .935$\pm$.007 .902$\pm$.001 .890$\pm$.008 .797$\pm$.001
CoDi N/A N/A .525$\pm$.021 .657$\pm$.049 .625$\pm$.013 .577$\pm$.054
STaSy N/A N/A .980$\pm$.001 .897$\pm$.002 .839$\pm$.006 .791$\pm$.001
TabSyn .916$\pm$.011 .690$\pm$.001 .984$\pm$.001 .895$\pm$.002 .839$\pm$.002 .789$\pm$.001
TRBD (ours) .643$\pm$.014 .619$\pm$.002 .922$\pm$.002 .901$\pm$.001 .977$\pm$.003 .801$\pm$.000
Table 14: Results of our experiments on the classification datasets. We use XGBoost as the classifier and F1 score as the metric.
CC AR BD AD CH CA
Real .620$\pm$.066 .595$\pm$.002 .927$\pm$.005 .689$\pm$.004 .585$\pm$.018 .719$\pm$.004
TVAE .$0$525$\pm$.003 .850$\pm$.012 .620$\pm$.008 .467$\pm$.014 .519$\pm$.017
CTGAN .$0$590$\pm$.005 .826$\pm$.002 .571$\pm$.030 .252$\pm$.026 .707$\pm$.005
TabDDPM .$0$353$\pm$.095 .573$\pm$.002 .661$\pm$.003 .669$\pm$.013 .724$\pm$.004
CoDi N/A N/A .333$\pm$.016 .409$\pm$.033 .154$\pm$.081 .650$\pm$.015
STaSy N/A N/A .841$\pm$.015 .614$\pm$.018 .592$\pm$.013 .727$\pm$.005
TabSyn .172$\pm$.066 .554$\pm$.003 .905$\pm$.005 .629$\pm$.014 .569$\pm$.002 .719$\pm$.003
TRBD (ours) .$0$485$\pm$.004 .568$\pm$.002 .663$\pm$.002 .854$\pm$.011 .724$\pm$.001
Table 15: Results of our experiments on the regression datasets. We use XGBoost as the classifier and $R^{2}$ score as the metric.
IS KI AB FB
Real .858$\pm$.021 .885$\pm$.003 .537$\pm$.042 .791$\pm$.007
TVAE .665$\pm$.049 .641$\pm$.006 .286$\pm$.007 .541$\pm$.043
CTGAN $-0.577$$\pm$.044 .580$\pm$.009 .288$\pm$.032 .420$\pm$.032
TabDDPM .903$\pm$.008 .817$\pm$.007 .553$\pm$.004 .615$\pm$.001
CoDi $-3.766$$\pm$.393 $-64.883$$\pm$14.632 $-1.583$$\pm$.322 N/A
STaSy .744$\pm$.039 .848$\pm$.011 $-0.201$$\pm$.075 .678$\pm$.016
TabSyn .843$\pm$.004 .835$\pm$.011 .515$\pm$.006 .638$\pm$.006
TRBD (ours) .915$\pm$.003 .900$\pm$.010 .568$\pm$.006 .689$\pm$.006
Table 16: Results of our experiments on the regression datasets. We use XGBoost as the classifier and RMSE as the metric.
IS KI AB FB
Real 4,777.171$\pm$690.880 129,298.734$\pm$8,184.660 2.301$\pm$.120 6.051$\pm$.098
TVAE 7,423.982$\pm$535.919 233,179.555$\pm$2,057.680 2.566$\pm$.013 9.084$\pm$.042
CTGAN 16,140.365$\pm$223.186 252,229.633$\pm$2,844.116 2.562$\pm$.057 10.405$\pm$.279
TabDDPM 3,872.860$\pm$156.814 166,357.334$\pm$2,973.649 2.031$\pm$.009 8.323$\pm$.135
CoDi 26,698.977$\pm$1,107.887 2,978,663.320$\pm$340,768.191 5.292$\pm$.346 N/A
STaSy 6,488.204$\pm$499.168 151,618.842$\pm$5,597.809 3.327$\pm$.103 7.612$\pm$.191
TabSyn 5,091.418$\pm$59.761 158,079.866$\pm$5,064.771 2.115$\pm$.013 8.079$\pm$.067
TRBD (ours) 3,630.228$\pm$69.255 123,291.742$\pm$5,934.827 1.995$\pm$.015 7.486$\pm$.069
Table 17: Results of our experiments on the classification datasets. We use RandomForest as the classifier and AUROC as the metric.
CC AR BD AD CH CA
Real .948$\pm$.011 .725$\pm$.001 .986$\pm$.002 .926$\pm$.002 .859$\pm$.004 .798$\pm$.004
TVAE Failed .633$\pm$.001 .980$\pm$.000 .876$\pm$.002 .763$\pm$.007 .718$\pm$.004
CTGAN Failed .648$\pm$.002 .965$\pm$.002 .883$\pm$.002 .680$\pm$.009 .766$\pm$.002
TabDDPM .511$\pm$.054 .499$\pm$.018 .978$\pm$.002 .906$\pm$.001 .895$\pm$.005 .799$\pm$.001
CoDi N/A N/A .551$\pm$.040 .601$\pm$.039 .630$\pm$.016 .562$\pm$.030
STaSy N/A N/A .980$\pm$.001 .898$\pm$.001 .836$\pm$.007 .794$\pm$.002
TabSyn .913$\pm$.007 .687$\pm$.001 .984$\pm$.001 .895$\pm$.002 .838$\pm$.002 .794$\pm$.001
TRBD (ours) .615$\pm$.054 .637$\pm$.002 .975$\pm$.002 .903$\pm$.001 .970$\pm$.003 .804$\pm$.001
Table 18: Results of our experiments on the classification datasets. We use RandomForest as the classifier and F1 as the metric.
CC AR BD AD CH CA
Real .597$\pm$.040 .572$\pm$.003 .930$\pm$.004 .689$\pm$.004 .527$\pm$.034 .716$\pm$.006
TVAE Failed .496$\pm$.005 .838$\pm$.001 .604$\pm$.013 .468$\pm$.030 .532$\pm$.017
CTGAN Failed .589$\pm$.003 .821$\pm$.005 .571$\pm$.030 .252$\pm$.026 .713$\pm$.002
TabDDPM .$0$136$\pm$.087 .790$\pm$.013 .661$\pm$.007 .671$\pm$.007 .717$\pm$.005
CoDi N/A N/A .321$\pm$.054 .395$\pm$.009 .067$\pm$.029 .630$\pm$.049
STaSy N/A N/A .821$\pm$.008 .599$\pm$.011 .583$\pm$.021 .728$\pm$.004
TabSyn .041$\pm$.027 .547$\pm$.001 .895$\pm$.006 .599$\pm$.019 .565$\pm$.021 .717$\pm$.003
TRBD (ours) .$0$479$\pm$.002 .751$\pm$.050 .661$\pm$.003 .864$\pm$.007 .728$\pm$.002
Table 19: Results of our experiments on the regression datasets. We use RandomForest as the classifier and $R^{2}$ score as the metric.
IS KI AB FB
Real .831$\pm$.052 .868$\pm$.009 .558$\pm$.032 .790$\pm$.007
TVAE .693$\pm$.031 .636$\pm$.010 .308$\pm$.006 .567$\pm$.056
CTGAN $-0.618$$\pm$.048 .569$\pm$.014 .281$\pm$.032 .436$\pm$.012
TabDDPM .903$\pm$.006 .817$\pm$.022 .547$\pm$.011 .653$\pm$.010
CoDi $-4.100$$\pm$.365 $-81.550$$\pm$10.043 $-2.019$$\pm$.417 N/A
STaSy .772$\pm$.018 .820$\pm$.011 $-0.500$$\pm$.046 .681$\pm$.013
TabSyn .861$\pm$.034 .825$\pm$.009 .515$\pm$.007 .656$\pm$.008
TRBD (ours) .913$\pm$.003 .899$\pm$.014 .559$\pm$.001 .695$\pm$.004
Table 20: Results of our experiments on the regression datasets. We use RandomForest as the classifier and RMSE as the metric.
IS KI AB FB
Real 4,737.314$\pm$330.580 137,629.411$\pm$7,872.392 2.106$\pm$.121 6.183$\pm$.141
TVAE 7,115.766$\pm$356.529 234,758.355$\pm$3,234.940 2.527$\pm$.011 8.819$\pm$.555
CTGAN 16,349.869$\pm$241.497 255,570.930$\pm$4,097.423 2.575$\pm$.058 10.086$\pm$.109
TabDDPM 3,880.795$\pm$115.701 166,210.395$\pm$9,867.884 2.044$\pm$.024 7.907$\pm$.111
CoDi 27,625.513$\pm$1,000.384 3,349,602.925$\pm$206,738.563 5.718$\pm$.406 N/A
STaSy 6,129.863$\pm$236.914 164,822.726$\pm$4,872.698 3.720$\pm$.058 7.581$\pm$.153
TabSyn 5,083.900$\pm$72.430 162,895.708$\pm$4,128.800 2.115$\pm$.014 7.876$\pm$.095
TRBD (ours) 3,676.466$\pm$66.751 131,291.742$\pm$7,757.758 2.017$\pm$.022 7.411$\pm$.054
D.2 Sample-wise Quality Score of Generated Data
We evaluate the generated data using statistical metrics. It is essential to note that these results should not be evaluated in isolation. As demonstrated in Section 4.3, accurate reproduction of categorical data does not inherently guarantee high quality. In other words, statistical excellence does not necessarily imply high quality. In generative tasks, the generated data is often used to train new machine learning models, and the primary evaluation should focus on downstream tasks. Hence, we consider statistical metrics as supplementary evaluations. It is not advisable to solely determine the superiority or inferiority of models based on these results. Following Liu et al. (2023) and Zhang et al. (2023), we use $\alpha$-Precision and $\beta$-Recall proposed in Alaa et al. (2022). $\alpha$-Precision evaluates how faithfully the generated data adheres to the distribution of real data. $\beta$-Recall assesses the coverage of the generated data.
Tables 21 and 22 present the result of the two metrics. TVAE and CTGAN achieve high scores in $\alpha$-Precision for the CC dataset, indicating a considerable fidelity to real data. However, as shown in Table 3, the generated data fails to reproduce the imbalance. This emphasizes that statistical metrics do not guarantee sampling quality.
In Sections 4.2, 4.3, and 4.4, TabSyn and TRBD demonstrated competitive results. This trend is also reflected in Table 21. However, as shown in Table 22, TRBD outperforms TabSyn in $\beta$-Recall. TabSyn transforms data into latent variables, potentially causing the loss of rare information during this process. Since the generative model cannot discern the importance of the lost rare information, the performance of TabSyn may degrade in more extensive experiments.
Table 21: Comparison of $\alpha$-Precision scores. Higher value indicates better score.
CC AR IS BD AD CH CA KI AB FB
TVAE .960$\pm$.001 .809$\pm$.001 .901$\pm$.009 .960$\pm$.004 .856$\pm$.004 .547$\pm$.005 .788$\pm$.002 .918$\pm$.002 .898$\pm$.011 .008$\pm$.002
CTGAN .861$\pm$.001 .915$\pm$.001 .722$\pm$.010 .739$\pm$.004 .717$\pm$.006 .894$\pm$.006 .746$\pm$.002 .465$\pm$.002 .594$\pm$.005 .934$\pm$.002
TabDDPM .$0$$0$974$\pm$.010 .973$\pm$.008 .933$\pm$.003 .933$\pm$.005 .975$\pm$.004 .968$\pm$.002 .908$\pm$.005 .578$\pm$.006
CoDi N/A N/A .368$\pm$.012 .105$\pm$.002 .031$\pm$.004 .387$\pm$.004 .070$\pm$.023 .033$\pm$.012 .059$\pm$.001 N/A
STaSy N/A N/A .359$\pm$.009 .871$\pm$.002 .638$\pm$.007 .718$\pm$.003 .933$\pm$.006 .830$\pm$.005 .829$\pm$.008 .902$\pm$.010
TabSyn .977$\pm$.001 .958$\pm$.001 .971$\pm$.010 .933$\pm$.002 .980$\pm$.004 .990$\pm$.004 .988$\pm$.002 .983$\pm$.007 .983$\pm$.007 .400$\pm$.005
TRBD (ours) .866$\pm$.001 .541$\pm$.001 .974$\pm$.008 .988$\pm$.007 .988$\pm$.003 .984$\pm$.007 .981$\pm$.002 .962$\pm$.003 .983$\pm$.006 .935$\pm$.002
Table 22: Comparison of $\beta$-Recall scores. Higher value indicates better score.
CC AR IS BD AD CH CA KI AB FB
TVAE .167$\pm$.001 .$0$274$\pm$.012 .201$\pm$.003 .347$\pm$.004 .090$\pm$.002 .265$\pm$.006 .194$\pm$.006 .179$\pm$.009 .000$\pm$.000
CTGAN .072$\pm$.001 .$0$159$\pm$.005 .271$\pm$.006 .267$\pm$.004 .232$\pm$.004 .330$\pm$.004 .110$\pm$.002 .008$\pm$.005 .086$\pm$.004
TabDDPM .$0$$0$694$\pm$.008 .565$\pm$.004 .503$\pm$.005 .506$\pm$.002 .502$\pm$.009 .376$\pm$.006 .458$\pm$.013 .261$\pm$.005
CoDi N/A N/A .110$\pm$.009 .006$\pm$.001 .$0$070$\pm$.003 .007$\pm$.001 .$0$002$\pm$.001 N/A
STaSy N/A N/A .063$\pm$.010 .326$\pm$.005 .306$\pm$.006 .365$\pm$.007 .471$\pm$.008 .350$\pm$.005 .440$\pm$.006 .394$\pm$.005
TabSyn .331$\pm$.001 .095$\pm$.001 .503$\pm$.014 .500$\pm$.005 .435$\pm$.006 .504$\pm$.003 .488$\pm$.007 .381$\pm$.006 .461$\pm$.006 .129$\pm$.005
TRBD (ours) .136$\pm$.001 .$0$742$\pm$.016 .498$\pm$.002 .445$\pm$.004 .650$\pm$.002 .499$\pm$.003 .548$\pm$.002 .485$\pm$.007 .182$\pm$.003
D.3 Runtime
We present the additional results of the runtime between TabDDPM and TRBD. The batch size for training is 4,096, and the number of iterations is kept the same for TabDDPM and TRBD. However, since the number of layers in the MLPs used in the architectures is different, we also provide an approximate count of the parameters.
Table 23 shows the comparison of the training time. From this table, it can be observed that even in cases where the number of parameters is higher, TRBD operates faster compared to TabDDPM, indicating that the computation of multinomial diffusion is the bottleneck. Additionally, on high cardinality datasets, despite having a higher number of parameters, poor results are obtained in terms of diversity and performance in downstream tasks. Therefore, it is unlikely that increasing the number of parameters will lead to a solution to the problem.
Table 23: Comparison of the training time and the number of parameters between TabDDPM and TRBD.
CC AR IS BD AD CH CA KI AB FB
Time TabDDPM 3,494s 3,014s 273s 471s 691s 567s 696s 600s 60.5s 648s
TRBD (ours) 183s 265s 113s 454s 495s 264s 404s 234s 41.9s 372s
#params TabDDPM 14.8M 13.1M 1.1M 1.3M 3.8M 1.3M 4.3M 5.4M 2.4M 0.7M
TRBD (ours) 0.4M 2.5M 1.5M 4.0M 4.9M 2.7M 3.7M 2.8M 1.6M 4.0M
Table 24 shows the comparison of the sampling time. We use the same model as during training and generate 300k samples for the CC and AR datasets, and 26k samples for the other datasets. The batch size during generation is set to match the number of samples generated, and the sampling steps are unified to $T=1000$.
It can be observed that similar results to those during training comparison are obtained. However, during generation, there is not a significant reduction in time, possibly because the bottleneck of multinomial diffusion is less prominent. Additionally, on high cardinality datasets, TabDDPM experiences OOM errors. This indicates that memory is exhausted due to the large dimensionality of one-hot vectors.
Table 24: Comparison of the sampling time and the number of parameters between TabDDPM and TRBD.
CC AR IS BD AD CH CA KI AB FB
Time TabDDPM OOM OOM 12.0s 14.0s 25.8s 22.8s 25.3s 14.5s 16.2s 21.4s
TRBD (ours) 4.3s 196s 9.4s 21.0s 24.3s 13.6s 18.2s 13.7s 11.3s 19.1s
#params TabDDPM 14.8M 13.1M 1.1M 1.3M 3.8M 1.3M 4.3M 5.4M 2.4M 0.7M
TRBD (ours) 0.4M 2.5M 1.5M 4.0M 4.9M 2.7M 3.7M 2.8M 1.6M 4.0M
D.4 ResBit with STaSy
We present the results of introducing ResBit to STaSy, along with the training times. For the evaluation metrics of TSTR, we use AUROC for the classification datasets and $R^{2}$ score for the regression datasets. We use CatBoost as the classifier.
Tables 25 and 26 present the results. It can be observed that training proceeded faster due to the reduction in input dimensions. Among the low cardinality datasets, AUROC is maintained for the classification datasets, and improvement in $R^{2}$ score is observed for most regression datasets. In high cardinality datasets, it can be confirmed that training completes within 24 hours due to the reduction in dimensions. For example, in the case of the AR dataset, with one-hot vectors, the dimensionality is $3+18+6571+292+293=7177$. However, using ResBit, it is reduced to fewer than $100$ dimensions.
Table 25: Results of TSTR framework in classification datasets. Metric is AUROC.
CC AR BD AD CH CA
TSTR STaSy N/A N/A .978$\pm$.002 .900$\pm$.002 .835$\pm$.004 .791$\pm$.001
STaSy with ResBit .785$\pm$.022 .549$\pm$.033 .980$\pm$.001 .898$\pm$.003 .834$\pm$.004 .790$\pm$.003
Training Time STaSy N/A N/A 4,707s 6,676s 3,879s 9,071s
STaSy with ResBit 57,854s 52,557s 3,211s 5,394s 2,408s 7,659s
Table 26: Results of TSTR framework in regression datasets. Metric is $R^{2}$ score.
IS KI AB FB
TSTR STaSy .618$\pm$.080 .867$\pm$.010 $-0.613$$\pm$.047 .670$\pm$.011
STaSy with ResBit .686$\pm$.007 .867$\pm$.007 .479$\pm$.005 .691$\pm$.010
Training Time STaSy 2,137s 4,937s 3,205s 26,406s
STaSy with ResBit 1,493s 3,477s 1,820s 24,494s
Appendix E Additional Applications of ResBit
We demonstrate the additional applications of ResBit. While it is applicable in scenarios involving one-hot vectors, we specifically explore its utility in image classification and class conditional image generation. It is important to emphasize that this demonstration does not claim an improvement in accuracy.
E.1 Image Classification
In the image classification model, we modify the output of the final layer. To ensure a fair comparison, we round output and evaluate its performance based on whether it matches the label representation (one-hot and ResBit). For one-hot vectors, a well-trained model should achieve performance comparable to top-1 accuracy. We use SGD (Sutskever et al., 2013) as the optimizer and Binary Cross Entropy with Sigmoid as the loss function. The dataset we use is Food 101 (Bossard et al., 2014). The learning rate is set as shown in Table 27, and ResNet34 (He et al., 2016) is employed as the classification model, with training conducted for 100 epochs with a batch size of 1,024.
Table 27: Learning rate in the training in the image classification.
Epochs (1-index) Learning Rate
1-30 0.1
31-60 0.02
61-80 0.004
81-100 0.0008
The results are presented in Table 28. The #Classes indicates the number of classes used from the dataset, with a total of 101 classes. While both methods show a decrease in accuracy with an increase in dimensionality, ResBit can mitigate the reduction in accuracy due to its ability to restrain dimensionality growth. However, in the case of one-hot vectors, the accuracy is only slightly higher than random guessing with 101 classes, indicating the difficulty of the task.
Table 28: Results of image classification experiments.
#Label dim Accuracy(%)
#Classes one-hot ResBit one-hot ResBit
10 10 5 52.32 50.00
20 20 7 34.38 34.04
40 40 9 20.47 26.67
60 60 16 9.97 18.15
80 80 11 6.31 24.85
101 101 15 1.61 21.59
E.2 Class Conditional Image Generation
We conduct experiments using Conditional Generative Adversarial Networks (CGAN) (Mirza & Osindero, 2014) as a method to specify classes. In other words, we concatenate a noise vector and a conditional vector representing the class and input them to the Generator. We use the CIFAR-10 (Krizhevsky, 2009), MNIST (LeCun et al., 2010), and Food101 datasets. For the architectures, we utilize InfoGAN (Chen et al., 2016) for CIFAR-10 and MNIST, and SNGAN (Miyato et al., 2018) for Food101. We use Adam (Kingma & Ba, 2017) as the optimizer with hyperparameters $lr=0.0002,\beta_{1}=0.5,\beta_{2}=0.999$. Our primary objective is not to improve the quality of generated images, so we do not incorporate accuracy improvement techniques. However, in the case of SNGAN, we aim for training stability by updating the discriminator five times for every single update of the generator. We conduct training for 30 epochs on MNIST, 300 epochs on CIFAR-10, and updated the generator 15,000 times for Food101. Results are presented in Figures 6, 7, and 8. It is evident that conditioning works for all datasets. In MNIST, ResBit lacks diversity, while in CIFAR-10, visually, it can be observed that one-hot has more mode collapse.
(a) One-hot conditioning
(b) ResBit conditioning (ours)
Figure 6: Class conditioned samples from InfoGAN trained on MNIST
(a) One-hot conditioning
(b) ResBit conditioning (ours)
Figure 7: Class conditioned samples from InfoGAN trained on CIFAR-10
(a) One-hot conditioning
(b) ResBit conditioning (ours)
Figure 8: Generated samples from SNGAN trained on Food101
For CIFAR-10, we conduct a comparison using the FID (Heusel et al., 2017) score. We calculate the FID score by generating 50,000 images for each method. The results are presented in Table 29. It is noticeable that there is no significant decrease in accuracy due to the change in the conditioning method. In addition, with conditioning, the “out-of-index” problem does not occur, making the conditioning with binary, which has the fewest dimensions, yield the best results.
In our experiments, we use a small number of classes. This decision is based on the observation that the FID score tends to degrade when the number of classes decreases or the number of images per class decreases in class-conditional GANs (Shahbazi et al., 2022).
Table 29: Comparison of the FID score on CIFAR-10 dataset.
one-hot ResBit (ours) Binary
FID 95.52 82.21 63.54
Label dim 10 5 4